OpenStudy (anonymous):

evaluate indefinite integral using the substitution method ..........again lol

OpenStudy (anonymous):

\[\int\limits_{}^{}x/\left( 1-x ^{^{2}} \right)^{-1/5}xdx\]

OpenStudy (anonymous):

sorry not -1/5 its 1/5

OpenStudy (anonymous):

i got \[-5/8\left( 1-x ^{2} \right)^{^{4/5}}+K\]

OpenStudy (anonymous):

is this right?

OpenStudy (experimentx):

Let \( 1 - x^2 = u\),then \( du = 2xdx\) \( =\large \frac{1}{2} \int \frac{2x}{(1 - x^2)^{-1/5}} dx\) \( =\large \frac{1}{2} \int \frac{du}{(u)^{-1/5}} dx\) \( =\large\frac{1}{2} [\frac{u^{5/6}}{5/6}] + C \) \( =\large\frac{3}{5} u^{5/6} + C \)

OpenStudy (experimentx):

\( 3/5 (1-x^2)^{5/6}\)

OpenStudy (anonymous):

wouldn't du=-2xdx?

OpenStudy (anonymous):

im sorry and its not ^(-1/5) its positive

OpenStudy (anonymous):

\[ \int \frac{x}{\sqrt[5]{1-x^2}} \, dx=-\frac{5}{8} \left(1-x^2\right)^{4/5} \] by following the method of @experimentX with the -1/5 replaced with 1/5

OpenStudy (anonymous):

thanks

OpenStudy (experimentx):

Sorry, there was mistake from second step \( = \large -\frac{1}{2} [\frac{u^{4/5}}{4/5}] + C\) \( = \large -\frac{5}{8} u^{4/5} + C\) \( = \large -\frac{5}{8} (1-x^2)^{4/5} + C\)

OpenStudy (anonymous):

omg thanks im so getting this now